a particle experiences constant acceleration for 6s after starting from rest. if it travels a distance s1 in the first 2sec and a distance s2 in the next 2 sec and a distance s3 in the last 2 sec . then s1:s2:s3 is equal to

Grade:11

1 Answers

Vikas TU
6 years ago
Dear Student,
Let the acceleration be a and the initial velocity=u=0

S1= ut+at2/2
=>S1=2a

velocity at the end of 2 sec = u+at = 2a

S2= ut+at2/2
=>S2=2a x 2 + ax4/2
=>S2=6a

Velocity at the end of 4 seconds= 4a
S3= ut+at2/2
=>S3=8a+4a/2
=>S3=10a

S1:S2:S3 = 1:3:5
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)